# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

### v3

@@ Abstract
Data-Independent Acquisition (DIA) was introduced to improve sensitivity to cover all peptides in a range rather than only sampling high-intensity peaks as in Data-Dependent Acquisition (DDA) mass spectrometry. However, it is not very clear how useful DIA data is for de novo peptide sequencing as the DIA data are marred with coeluted peptides, high noises, and varying data quality. We present a new deep learning method DIANovo, and address each of these difficulties, and improves the previous established system DeepNovo-DIA by from 25% to 81%, averaging 48%, for amino acid recall, and by from 27% to 89%, averaging 57%, for peptide recall, by equipping the model with a deeper understanding of coeluted DIA spectra. This paper also provides criteria about when DIA data could be used for de novo peptide sequencing and when not to by providing a comparison between DDA and DIA, in both de novo and database search mode. We find that while DIA excels with narrow isolation windows on older-generation instruments, it loses its advantage with wider windows. However, with Orbitrap Astral, DIA consistently outperforms DDA due to narrow window mode enabled. We also provide a theoretical explanation of this phenomenon, emphasizing the critical role of the signal-to-noise profile in the successful application of de novo sequencing.
@@ 1 Introduction
Recent advancements such as DeepNovo-DIA [19] and Transformer-DIA [20] have harnessed the power of deep learning to address the challenges inherent to de novo sequencing in DIA. These models employ neural networks to analyze precursor and fragment ions across multiple dimensions, including mass-to-charge ratio (m/z), retention time, and intensity, thus effectively handling the complexity of highly multiplexed spectra. For instance, DeepNovo-DIA integrates Ion-CNN and Spectrum-CNN  [21] with a long short-term memory (LSTM)  [22] network to capture the three-dimensional structure of fragment ions and their interrelationships. Similarly, Transformer-DIA adopts a Transformer-based  [23] architecture, using an encoder-decoder framework to predict peptide sequences by iteratively generating each subsequent amino acid based on prior outputs. Despite their efficacy, both models are constrained by their focus on encoding only a portion of the spectrum, potentially missing important signals related to future amino acids.
@@
Our extensive experiments across various datasets, including older-generation and Astral data, highlight the robustness and effectiveness of our proposed method. We demonstrated that our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA. Sensitivity analysis further reveals that our algorithm maintains stable performance even with any number of coeluting peptides. These findings underscore the potential of our method for robust and accurate de novo peptide sequencing in the challenging DIA setting.
@@ 2 Results
Our experiments comprehensively evaluated the performance and robustness of our de novo peptide sequencing algorithm across various datasets and conditions. We found that older-generation mass spectrometers exhibit relatively lower peptide recall, while Astral data showed better results due to improved fragment ion coverage. Our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA model across multiple datasets, highlighting its superior capability in complex DIA settings.
We also assessed the algorithm’s sensitivity to coelution numbers (number of coeluting peptides), revealing its robust performance even with high numbers of coeluting peptides, as evidenced by the sensitivity study with Yeast KO [30] dataset with an average coelution number of 19. These coeluting peptides, characterized by overlapping retention times and shared precursor windows, pose significant challenges due to spectral complexity. Our model outperforms the baselines, by better distinguishing target peptides from coeluting ones. This improvement stems from features like coelution-aware pretraining and robust spectrum encoding, enabling more accurate peptide identification in complex DIA data.
@@ 2.1 De Novo Performance on Older-Generation Data
We compared our model’s performance with the baselines across several datasets [34] [35] [33] in Figure 1. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board, indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 53% higher than DeepNovo-DIA, while peptide recall being 56% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 27% and 32% higher respectively.
@@
Figure 1: Amino acid recall (a) and peptide recall (b) of our method vs DeepNovo-DIA and Transformer-DIA, training sequences excluded from test set, on various older-generation datasets, measured over 10,000 randomly selected peptide precursors per dataset.
@@ 2.2 Performance on Orbitrap Astral Data
Figure 2: Amino acid recall (a) and peptide recall (b) of our method vs baselines on various Astral datasets.
The performance of our model on the Astral datasets shows a marked improvement over the baselines. On average we can expect a 45%/57% increase of amino acid / peptide recall with our methods compared to DeepNovo-DIA, or a 22%/38% improvement compared to Transformer-DIA. Furthermore, Astral data delivers a significant boost to de novo performance, affirming the effectiveness of DIA de novo sequencing in such system.
2.3 Sensitivity to Coelution Number
To visualize the relationship between coelution number and de novo sequencing performance, we generated a plot (shown in Figure 3) that demonstrates our algorithm’s performance across different levels of coelution. The plot illustrates that our algorithm maintains consistent performance irrespective of the coelution number, indicating its robustness in handling complex spectral data. Compared to DeepNovo-DIA and Transformer-DIA, which experiences a drop in performance as coelution increases, our method demonstrates a significant stability advantage. This robustness is achieved through the careful design of our algorithm to address the complexities inherent in coeluted data. By leveraging coelution-aware pretraining and advanced spectrum encoding techniques, such as dilated CNNs and self-attention mechanisms, our model effectively distinguishes target peptide signals from overlapping fragment ions of coeluting peptides.
((a)) Amino acid recall
((b)) Peptide recall
Figure 3: Amino acid (a) and peptide (b) recall of our method vs baselines, on the Yeast KO Dataset, on peptides with varying coelution number, measured on a total of 20,000 randomly selected peptide precursors.
@@ 2.5 Ablation Study
In this ablation study, illustrated in Figure 6, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The results highlight that coelution-aware pretraining plays a crucial role in model performance, as its removal reduces peptide recall from 31.2% to 28.1%. More notably, the removal of Rotary Positional Encoding (RoPE) leads to a severe performance drop, with peptide recall plunging to 21.2%. This decline suggests that absolute positional embeddings alone are insufficient for encoding graph node masses effectively. Additionally, excluding extended peak features or limiting the input to only three neighboring spectra results in moderate performance losses, demonstrating the importance of both extended features and broader spectral context. Finally, binning spectra into 0.5 Th intervals causes a significant drop in performance due to reduced spectral resolution, which increases ambiguity in peptide identification. The substantial degradation observed in these ablations underscores the essential role of these components in enhancing the model’s ability to identify peptides accurately.
((a)) Amino acid recall
((b)) Peptide recall
Figure 6: Amino acid (a) and peptide (b) recall of our method vs different configurations, on the Yeast KO dataset.
@@ 2.6 Theoretical Analysis of De Novo Peptide Sequencing Performance
    Figure 7: Simulated p-values vs peptide recall for test datasets, with a Pearson correlation coefficient of -0.58.
@@ 3 Discussion
Our comprehensive experiments demonstrate the superior performance of the proposed method, particularly when applied to data from the Orbitrap Astral mass spectrometer. The Astral’s enhanced fragment ion coverage and consistent fragmentation patterns provide a solid foundation for our de novo sequencing approach, leading to a marked increase in peptide identification accuracy, even in case of high coelution level. In contrast, older-generation mass spectrometers, such as those used in older DDA and DIA experiments, reveal the limitations of existing de novo sequencing methodologies, especially when larger isolation windows are used. Our DDA vs. DIA comparison study highlights these challenges, showing that while DIA outperforms DDA with narrower isolation windows, it begins to lose this advantage as the window size increases. This issue is particularly evident with older-generation instruments, where DIA performance in both database search and de novo sequencing diminishes at larger window sizes. However, the Orbitrap Astral consistently demonstrates DIA’s superiority over DDA, facilitated by the narrow isolation windows enabled by this instrument, underscoring the pivotal role that modern mass spectrometers play in advancing peptide sequencing capabilities.
Furthermore, the sensitivity analysis reinforces the robustness of our method, showing stable performance even in scenarios with high numbers of coeluting peptides—common in proteomics workflows. Maintaining high peptide recall under these challenging conditions is essential for practical applications in the field, where complex mixtures and high levels of coelution are routine. Our method’s ability to handle these complexities underscores its versatility and reliability across a range of proteomic datasets. In addition, our ablation study verifies the validity and neccessity of our design.
@@ 4.4 Two Stage Decoder
Reader can refer to Figure 1a of Mao et al. (2023)[13] for a detailed visualization of the two-stage decoding process.
@@ Appendix E Characteristics of Test Datasets
List of supportingtables 3 Experimental Parameters for older-generation test datasets. Coeluting number refers to how many coeluting peptides one peptide has on average, number of signal or noise peaks refers to the median number of peaks of the neighboring five spectrums which are fragment ions or the target peptide or not, median noise intensity refers to the ratio between median intensity for noise peaks and signal peaks, and isolation window has unit Th.
@@ Appendix F Links to Datasets
Supporting Table 4 provides links to our training and testing datasets.
@@ Appendix G Model Implementation Details
Our model is trained on a Lion optimizer [47] with learning rate $2\times 10^{-6}$, over 10 epochs. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453, with 1,896,662 and 1,086,577 PSMs respectively.

### v4

@@ Abstract
Data-Independent Acquisition (DIA) was introduced to improve sensitivity to cover all peptides in a range rather than only sampling high-intensity peaks as in Data-Dependent Acquisition (DDA) mass spectrometry. However, it is not very clear how useful DIA data is for de novo peptide sequencing as the DIA data are marred with coeluted peptides, high noises, and varying data quality. We present a new deep learning method DIANovo, and address each of these difficulties, and improves the previous established system DeepNovo-DIA by from 34% to 108%, averaging 50%, for amino acid recall, and by from 32% to 83%, averaging 57%, for peptide recall, by equipping the model with a deeper understanding of coeluted DIA spectra. This paper also provides criteria about when DIA data could be used for de novo peptide sequencing and when not to by providing a comparison between DDA and DIA, in both de novo and database search mode. We find that while DIA excels with narrow isolation windows on older-generation instruments, it loses its advantage with wider windows. However, with Orbitrap Astral, DIA consistently outperforms DDA due to narrow window mode enabled. We also provide a theoretical explanation of this phenomenon, emphasizing the critical role of the signal-to-noise profile in the successful application of de novo sequencing.
@@ 1 Introduction
Recent advancements such as DeepNovo-DIA [19], Transformer-DIA [20], and Cascadia [21] have harnessed the power of deep learning to address the challenges inherent to de novo sequencing in DIA, while approaches such as PepNet [22] have sought to develop unified models capable of handling both DDA and DIA data modalities.. These models employ neural networks to analyze precursor and fragment ions across multiple dimensions, including mass-to-charge ratio (m/z), retention time, and intensity, thus effectively handling the complexity of highly multiplexed spectra. For instance, DeepNovo-DIA integrates Ion-CNN and Spectrum-CNN  [23] with a long short-term memory (LSTM)  [24] network to capture the three-dimensional structure of fragment ions and their interrelationships. Similarly, Transformer-DIA adopts a Transformer-based  [25] architecture, using an encoder-decoder framework to predict peptide sequences by iteratively generating each subsequent amino acid based on prior outputs. Despite their efficacy, both models are constrained by their focus on encoding only a portion of the spectrum, potentially missing important signals related to future amino acids.
@@
Our extensive experiments across various datasets, including older-generation and Astral data, highlight the robustness and effectiveness of our proposed method. We demonstrated that our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA. These findings underscore the potential of our method for robust and accurate de novo peptide sequencing in the challenging DIA setting.
@@ 2 Results
Our experiments comprehensively evaluated the performance and robustness of our de novo peptide sequencing algorithm across various datasets and conditions. We found that older-generation mass spectrometers exhibit relatively lower peptide recall, while Astral data showed better results due to improved fragment ion coverage. Our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA model across multiple datasets, highlighting its superior capability in complex DIA settings. We also include a comparison with a recent state-of-the-art model, Cascadia, demonstrating the superior performance of our approach.
@@ 2.1 De Novo Performance on Older-Generation Data
We compared our model’s performance with the baselines across several datasets [36] [37] [35] in Figure 1, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board, indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
@@
((c)) Amino acid precision
Figure 1: Amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs DeepNovo-DIA and Transformer-DIA, training sequences excluded from test set, on various older-generation datasets, measured over 10,000 randomly selected peptide precursors per dataset.
@@ 2.2 Performance on Orbitrap Astral Data
((c)) Amino acid precision
Figure 2: Amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs baselines on various Astral datasets.
The performance of our model on the Astral datasets shows a marked improvement over the baselines. On average we can expect a 43%/60% increase of amino acid / peptide recall with our methods compared to DeepNovo-DIA, or a 27%/47% improvement compared to Transformer-DIA. Furthermore, Astral data delivers a significant boost to de novo performance, affirming the effectiveness of DIA de novo sequencing in such system.
@@ 2.3 Comparison of Peptide Detection by DDA and DIA
2.4 Comparison with Cascadia [21]
To evaluate the performance of our approach in a practical setting, we conduct a comparative analysis against Cascadia, a recently introduced state-of-the-art model. For this purpose, we selected a representative raw file from the PXD046386 Yeast Knockout (KO) dataset. This dataset provides a relevant benchmark for assessing model effectiveness in proteomics data analysis. Both our method and Cascadia were applied to this identical data source under equivalent preprocessing and parameter settings to ensure a fair comparison. The resulting performance differences are summarized and visualized in Figure 5.
Figure 5: Venn diagram for Cascadia comparison.
Our model demonstrates a 21% increase in peptide identifications compared to Cascadia. Notably, the overlap in identified peptides between the two models is relatively limited, indicating that each model captures distinct subsets of the data. This suggests that our approach exhibits different identification characteristics, potentially offering complementary insights to those provided by Cascadia.
@@ 2.5 Ablation Study
In this ablation study, illustrated in Figure 6, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The results highlight that coelution-aware pretraining plays a helpful role in model performance, as its removal reduces peptide recall by 8.5% compared to DIANovo. More notably, the removal of Rotary Positional Encoding (RoPE) leads to a severe performance drop, with peptide recall plunging to 67.2% of the original. This decline suggests that absolute positional embeddings alone are insufficient for encoding graph node masses effectively. Additionally, excluding extended peak features or limiting the input to only three neighboring spectra results in moderate performance losses, demonstrating the importance of both extended features and broader spectral context. Finally, binning spectra into 0.5 Th intervals causes a significant drop in performance due to reduced spectral resolution, which increases ambiguity in peptide identification. We also included a version of our model without binning the spectrum. In this setting, we adopt a Casanovo[12]-style approach where raw peaks are encoded individually, and the retention time (RT) dimension is represented using sinusoidal positional embeddings. Our results show that although the no-bin version reduces preprocessing time, the binned model outperforms the no-bin variant significantly. We believe this improvement stems from the fact that the binned approach enables the model to better utilize the retention time dimension through its more structured and compact encoding. The degradation observed in these ablations underscores the essential role of these components in enhancing the model’s ability to identify peptides accurately.
((a)) Relative amino acid recall
((b)) Relative peptide recall
((c)) Relative amino acid precision
Figure 6: Relative amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs different configurations, compared to our method, on the Yeast KO dataset.
@@ 2.6 Theoretical Analysis of De Novo Peptide Sequencing Performance
    Figure 7: Simulated p-values vs peptide recall for test datasets, with a Pearson correlation coefficient of -0.68.
@@ 3 Discussion
Our comprehensive experiments demonstrate the superior performance of the proposed method, particularly when applied to data from the Orbitrap Astral mass spectrometer. The Astral’s enhanced fragment ion coverage and consistent fragmentation patterns provide a solid foundation for our de novo sequencing approach, leading to a marked increase in peptide identification accuracy, even in case of high coelution level. In contrast, older-generation mass spectrometers, such as those used in older DDA and DIA experiments, reveal the limitations of existing de novo sequencing methodologies, especially when larger isolation windows are used. Our DDA vs. DIA comparison study highlights these challenges, showing that while DIA outperforms DDA with narrower isolation windows, it begins to lose this advantage as the window size increases. This issue is particularly evident with older-generation instruments, where DIA performance in both database search and de novo sequencing diminishes at larger window sizes. However, the Orbitrap Astral consistently demonstrates DIA’s superiority over DDA, facilitated by the narrow isolation windows enabled by this instrument, underscoring the pivotal role that modern mass spectrometers play in advancing peptide sequencing capabilities. In addition, our ablation study verifies the validity and neccessity of our design.
@@ 4.4 Two Stage Decoder
Readers can refer to Figure 1a of Mao et al. (2023)[13] for a detailed visualization of the two-stage decoding process.
@@ Appendix E Characteristics of Test Datasets
List of supportingtables 3 Experimental Parameters for test datasets. Coeluting number refers to how many coeluting peptides one peptide has on average, number of signal or noise peaks refers to the median number of peaks of the neighboring five spectrums which are fragment ions or the target peptide or not, median noise intensity refers to the ratio between median intensity for noise peaks and signal peaks, and isolation window has unit Th.
@@ Appendix F Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 [29] with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 [46].
@@ Appendix G Model Implementation Details
Our model is trained on a Lion optimizer [49] with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453, with 1,896,662 and 1,086,577 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
We report the training and validation metrics for three datasets: the older-generation dataset, PXD046453, and PXD046386. The results are broken down by model component.
Older-Generation Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.532 0.399
    Ion Accuracy (%) 82.8% 86.8%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000134 0.000114
    Node Recall (%) 84.4% 87.3%
    Node Precision (%) 93.7% 94.7%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.337 0.335
PXD046453 Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.422 0.312
    Ion Accuracy (%) 92.6% 92.6%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000171 0.000124
    Node Recall (%) 74.8% 81.0%
    Node Precision (%) 89.9% 92.8%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.537 0.535
PXD046386 Dataset
  • Pretraining Model
    Metric Training Validation
    Cross-Entropy Loss 0.414 0.412
    Ion Accuracy (%) 87.2% 87.6%
  • Optimal Path Model
    Metric Training Validation
    Focal Loss 0.000100 0.0000944
    Node Recall (%) 87.4% 89.1%
    Node Precision (%) 93.9% 93.4%
  • Sequence Generation Model
    Metric Training Validation
    Cross-Entropy Loss 0.130 0.171
To prevent information leakage and ensure a rigorous evaluation of model generalization, we designed our training/validation/test splits to fully separate related peptide signals across sets, including coeluting peptides. Our model architecture leverages not only the target peptide but also contextual information from coeluting peptides within a spectrum. Therefore, to eliminate any direct or indirect sequence overlap, we adopted a raw file–based splitting strategy: for each training run, we designated one raw file exclusively for validation and used the remaining files for training. This guarantees that no peptides, or their coeluting partners, appear in both training and validation sets.
Each training scheme was associated with a held-out validation set comprising 20,000 peptide-spectrum matches (PSMs). These validation examples were used strictly for hyperparameter tuning and early stopping. No validation data was used in model selection or test-time inference.
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. Furthermore, in our cross-species evaluation using Orbitrap Astral data, we imposed an even stricter criterion: training and testing data come from different species. Specifically, when training on the human dataset PXD046453, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing. This species-based isolation helps ensure that learned representations do not exploit taxon-specific sequence features or instrument-specific noise patterns.
This design aligns with the DOME [50] guidelines by ensuring that spectra from the same peptide, protein, or coeluting group are never split across training and test partitions. All splits are reproducible and deterministic based on raw file boundaries.
@@ Appendix I Sensitivity Analysis
Appendix J Attention Pattern Analysis between RoPE and Absolute Positional Embeddings
To understand how positional encoding methods influence the modeling of fragment ions in de novo peptide sequencing, we compare rotary positional embeddings (RoPE) with absolute positional embeddings. Specifically, we examine how each method affects the model’s attention patterns and hidden representations of fragment ions.
Our analysis focuses on two aspects: attention scores between fragment ion pairs and cosine similarity between their hidden embeddings. Fragment ion pairs are categorized into four groups: ions from the same cleavage site, from different cleavage sites, one amino acid apart, and from distant amino acids.
As shown in Supporting Tables 5 and 6, RoPE consistently produces higher attention scores and embedding similarities across all categories. For ions from the same cleavage site, the attention score under RoPE is 0.166, compared to 0.119 with absolute PE. Similar trends hold for embedding similarity, where ions one amino acid apart reach 0.801 under RoPE, versus 0.523 with absolute PE. Even for distant ions, RoPE maintains stronger relationships, indicating a better ability to capture long-range dependencies.
These results suggest that RoPE encourages the model to focus more on functionally related ions and to represent them more coherently. This improved relational modeling likely enhances the model’s ability to interpret peptide fragmentation patterns effectively.
Category RoPE Attention Score Absolute PE Attention Score
Fragment ions from same cleavage site 0.119 0.166
Fragment ions from different cleavage sites 0.015 0.007
Fragment ions from one amino acid apart 0.054 0.010
Fragment ions from distant amino acids 0.002 0.001
List of supportingtables 5 Attention scores for different types of ions.
Category RoPE Embedding Similarity Absolute PE Emebdding Similarity
Fragment ions from same cleavage site 0.885 0.826
Fragment ions from different cleavage sites 0.719 0.454
Fragment ions from one amino acid apart 0.801 0.523
Fragment ions from distant amino acids 0.688 0.218
List of supportingtables 6 Embedding cosine similarities for different types of ions.
Appendix K Sensitivity to Coelution Number
To visualize the relationship between coelution number and de novo sequencing performance, we generated a plot (shown in Supporting Figure 5) that demonstrates our algorithm’s performance across different levels of coelution. The plot illustrates that our algorithm maintains consistent performance irrespective of the coelution number, indicating its robustness in handling complex spectral data.
List of supportingfigures 5 Performance of our method vs baselines, on the Yeast KO Dataset, on peptides with varying coelution number.